OpenStudy (anonymous):

Integrated Interval questions? Please Help! 1.10/(square root of x)+10(square root of x) 2.(x^3)/3+5x^4

myininaya (myininaya):

what is the question?

myininaya (myininaya):

\[\int\limits\limits_{}^{} (\frac{10}{\sqrt{x}}+10\sqrt{x}) dx\]

OpenStudy (anonymous):

How to evaluate the indefinite integral of the first equation? And how to find the indefinite integral of the second. I'm sorry I was unclear.

myininaya (myininaya):

is the way i wrote it right?

OpenStudy (dumbcow):

use this general rule \[\int\limits x^n = \frac{x^{n+1}}{n+1} +C\]

OpenStudy (anonymous):

Yeah its the way you wrote it!

myininaya (myininaya):

ok use what dumbcow said you might want to recall sqrt(x)=x^(1/2)

OpenStudy (anonymous):

I'm not sure I understand how to apply this formula.. For the first one I got 20sqrt{x}+20x^3/2/3+c. Do you know if that would be correct?

myininaya (myininaya):

looks good